# Multiband Gravitational Wave Detection Prospects for M31 UCXB-1 System in Low and Middle Frequency Band

### v1

@@ Abstract
The recent discovery of M31 UCXB-1, the first extragalactic ultracompact X-ray binary (UCXB) with an orbital period of $T_{\rm orb}\sim 465$ s, presents a unique laboratory for studying close binary evolution and an unprecedented target for continuous gravitational wave (GW) searches. Its identification as a strong candidate black hole-white dwarf (BH-WD) system, combined with its exceptionally short period and high X-ray luminosity, suggests it may be one of the most vital low-frequency GW sources in M31. In this Letter, we investigate the detectability of its GW signal for future space-borne detectors in multiband GW detection. We find that while its signal-to-noise ratio (S/N) for low-frequency detectors remains marginal for high-confidence detection, middle-frequency detectors such as DECIGO and BBO are far more promising, potentially achieving S/N $\varrho>8$ within reasonable observational duration. With a primary mass of only $m_{1}>5.4M_{\odot}$ (or $6.6M_{\odot}$), the network of all low and middle frequency detector (or BBO alone) is capable of detecting GW from this system with a $\varrho>8$, during 10-year observation. Furthermore, orbital eccentricity can enhance the GW strain at higher harmonics, further improving its detectability, especially for middle-frequency detectors. This study establishes M31 UCXB-1 as a key prototype of short-period UCXBs, cementing its role as a cornerstone for multiband, multi-messenger astrophysics and a vital bridge between X-ray astronomy and the future GW era.
@@ I Introduction
Ultracompact X-ray binaries (UCXBs), with orbital periods below $\sim 80$ minutes, represent the end state of low-mass X-ray binary evolution involving a compact accretor [neutron star (NS) or black hole (BH)] and a hydrogen-deficient donor (2021MNRAS.506.4654W; 2023pbse.book.....T). They are powerful laboratories for studying extreme accretion physics, tidal interactions, and angular momentum loss mechanisms (e.g., 2016ApJ...830..131C; 2021MNRAS.503.3540C). Furthermore, their tight orbits make them prominent sources of persistent, low-frequency gravitational waves (GWs) driven by radiation reaction, placing them squarely within the sensitivity bands of proposed space-borne interferometers (2009CQGra..26i4030N; 2020ApJ...900L...8C; 2020ApJ...896..129C; 2023ApJ...944...83Q; 2021MNRAS.503.5495S; 2023ApJ...953..153H; 2025ApJ...995...27X). including Laser Interferometer Space Antenna (LISA, 2017arXiv170200786A), Tianqin (2016CQGra..33c5010L), and Taiji (2017SSPMA..47a0404H).
@@
Recently, a remarkable system has emerged from deep X-ray monitoring of the Andromeda galaxy (M31). 2024MNRAS.530.2096Z reported the discovery of a periodic X-ray source (RX J0042.3+4115/3XMM J004222.9+411535) in the M31 bulge with a period of $T_{\rm orb}\approx 465$ s and a peak luminosity of $L_{X}\sim 2.9\times 10^{38}$ erg s-1. Follow-up spectral analysis by 2025arXiv251202360M strongly argues that this source, designated M31 UCXB-1, is the first known extragalactic UCXB and a compelling BH-white dwarf (WD) candidate. It possesses the shortest orbital period among known UCXBs. The unique parameters of M31 UCXB-1—a likely BH-WD binary with a $\sim 10^{2}$-s period at a distance of 785 kpc—make its GW emission exceptionally interesting. The expected GW frequency ($f_{\rm gw}\approx 4.3$ mHz) falls within the sensitivity range of low-frequency GW observatories. Previous estimates suggest it may be detectable by LISA/Taiji if the BH mass is sufficiently high (2025ApJ...986..219Y; 2025arXiv251202360M). However, a detailed and systematic assessment of its detectability across the full landscape of future GW missions, considering middle-frequency detectors, mission durations, parameter estimation error and the potential impact of orbital eccentricity, is still lacking.
@@
In this Letter, we perform a comprehensive investigation into the GW detectability of M31 UCXB-1 in low and middle frequency band. We go beyond simple characteristic strain comparisons and compute expected signal-to-noise ratios (S/N) for a suite of planned low and middle frequency detectors, including LISA, Taiji, TianQin, DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) (2006CQGra..23S.125K), and Big Bang Observer (BBO) (2006CQGra..23.4887H). We explore a realistic range of system parameters (BH mass, orbital eccentricity) informed by the latest observational and theoretical constraints (2024MNRAS.530.2096Z; 2025arXiv251202360M; 2025ApJ...986..219Y). Our analysis reveals that middle-frequency detectors offer a superior pathway for a high-confidence detection of this source. Furthermore, we quantify that if the system exhibits non-zero orbital eccentricity, it can enhance the detection S/N for both low- and middle-frequency GW detectors, with a particularly pronounced improvement for middle-frequency detectors. This work solidifies M31 UCXB-1’s role as a prime multimessenger and multiband target and provides critical predictions for the observing strategies of forthcoming GW observatories in low and middle frequency band.
@@
This letter is structured as follows: In Section II, we assess the GW detectability of M31 UCXB-1 and the associated parameter estimation uncertainties for low- and middle-frequency GW detectors, under the assumption of a circular orbit. Subsequently, in Section III, we extend the analysis to include eccentric orbital configurations. Finally, we summarize our findings and present conclusions in Section IV.
@@ II.1 Parameter Setting
The binary system is located at a distance of approximately 785 kpc from Earth. Its orbital period is about 465 s, corresponding to an orbital frequency of $2.15\times 10^{-3}$ Hz. The WD in the system has a mass of $0.08761\,M_{\odot}$, while the mass of the black hole remains uncertain due to its degeneracy with the orbital inclination in X-ray observations (2025arXiv251202360M). A BH mass of $m_{1}=3M_{\odot}$ corresponds to an inclination $\iota\geq 73.5^{\circ}$, while $m_{1}=20M_{\odot}$ corresponds to $\iota\geq 86.7^{\circ}$ (2025arXiv251202360M). Since the amplitude of GW is not sensitive to the inclination $\iota$ (see Table 2), and eclipsing phenomena reflects this system is basically edge on. Unless otherwise specified, throughout this work we adopt an edge-on orbital inclination of $\iota=90^{\circ}$. Some fixed parameters about the system are summarized in Table 1. Since 2025arXiv251202360M assume that its orbit is circular, we just consider circular orbit in this Section, and we will consider its eccentric orbit in next Section (Sec. III).
@@ II.2 GW Signals
The GW radiated from a binary system can be expressed as (maggiore2008gravitational; 2025ApJ...978..104G)
@@
where $h_{0}=\frac{4{G}^{5/3}\mathcal{M}^{5/3}(\pi f)^{2/3}}{{c}^{4}d}$, $f=2f_{\rm orb}$ is the frequency of GW, $A_{+}=\frac{1+\cos^{2}\iota}{2}$, $A_{\times}=\cos\iota$, $\times$ represent different polarizations, $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass of this system and $\Phi(t)$ is the phase of GW. The root-mean-square (RMS) average of GW strain on all polarization is h_rms=⟨h_+^2+h_×^2⟩=h_0A+2+A×22. Adopting matched filtering method, the characteristic strain for GW is (see 2015CQGra..32a5014M; 2005ApJ...623...23S)
@@
where $N_{\rm eff}=\min\left\{\frac{f^{2}}{\dot{f}},T_{\rm obs}f\right\}$(2022ApJ...939...55G; 2025ApJ...978..104G), the variation rate of frequency is $\dot{f}=\frac{96}{5}\pi^{8/3}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{5/3}f^{11/3}$. For this binary system, the relation $\frac{f^{2}}{\dot{f}}\gg T_{\mathrm{obs}}f$ holds, implying that the evolution timescale of the GW frequency is much longer than the coherent observing duration. Consequently, the characteristic strain $h_{\mathrm{c}}$ depends primarily on the observation duration $T_{\mathrm{obs}}$.
@@
And its merger time scale due to GW radiation can be estimated by (maggiore2008gravitational)
@@
Therefore, we can summarize several typical characteristic strains and other dynamical parameters for different cases in Table 2. Although the accretion process of a black hole can alter the masses $m_{1}$ and $m_{2}$ of the binary system, the resulting mass change is negligible over the observational timescale (<10<10 yr) due to low accretion rate $\dot{M}\sim 10^{-8}M_{\odot}$/yr (2025ApJ...986..219Y)]. Therefore, in our study on GWs, we neglect the mass variation due to accretion as well as the potential changes in orbital dynamics.
@@ II.3 GW Detectors
Tianqin $10^{-4}$ 0.5 2020PhRvD.101j3027L
LISA $10^{-4}$ 0.5 2019CQGra..36j5011R
Taiji $10^{-4}$ 0.5 2023PhRvD.107f4021L
DECIGO $10^{-3}$ $10^{2}$ 2011PhRvD..83d4011Y; 2017PhRvD..95j9901Y
BBO $10^{-3}$ $10^{2}$ 2011PhRvD..83d4011Y; 2017PhRvD..95j9901Y
It is likely that this system will be detected by future low-frequency and middle-frequency GW detectors. The low-frequency GW detectors encompass LISA, Taiji, and Tianqin, whereas the middle-frequency GW detectors include DECIGO and BBO in this paper. For all GW detectors used in this work, their noise power spectral density $S_{\rm n}(f)$ are summarized in the references provided in Table 3 and their detectable frequency ranges are also summarized in Table 3. We show some typical characteristic strain values of this GW source (red and green points/diamonds, Eq. 4) and noise curve of all kinds of GW detectors in low and middle frequency band in the Figure 1. The characteristic strain values computed by us are broadly consistent with those reported by 2025ApJ...986..219Y; 2025arXiv251202360M. The slightly lower amplitude in our results is due to our choice of orbital inclination $\iota=90^{\circ}$.
@@
Figure 1: Possible characteristic strain $h_{\rm c}(f)$ of GW emitted from this system [red ($m_{1}=3M_{\odot}$) and green ($m_{1}=20M_{\odot}$) points($T_{\rm obs}=4$ yr)/diamond($T_{\rm obs}=10$ yr)] and noise curve $h_{\rm n}(f)$ for different GW detectors in low and middle frequency band with different colors represent different GW detectors as shown in the legend. These dotted lines for Taiji and LISA represent the noise curves of GW detectors without Galactic binaries foreground confusion (See 2023PhRvD.107f4021L). And this black line with triangle (red line with $\times$) symbols represents characteristic strain $h_{{\rm c},i}$ for eccentric orbit with e=0.2e=0.2(0.4), $m_{1}=3M_{\odot}$, and $T_{\rm obs}=4$ yr.
@@ II.4 S/N
The S/N of GW from this system detected by a GW detector is (2015CQGra..32a5014M; 2015CQGra..32e5004M)
@@
For a network of GW detectors, its effective S/N is calculated as follows(2022ApJ...939...55G; 2025ApJ...978..104G)
@@
where $\varrho_{i}$ is the S/N for each detector $i$, the sum is taken over the $N_{\rm d}$ detectors. In this letter, we mainly consider two GW detector networks: Tianqin-LISA-Taiji low frequency GW detector network and Tianqin-LISA-Taiji-DECIGO-BBO (TLTDB) network including all low and middle frequency GW detectors in this letter. Through calculation, we obtain its S/Ns for different detectors as a function of $m_{1}$, as shown in Figure 2 ($T_{\rm obs}=4$ yr) and Figure 3 ($T_{\rm obs}=10$ yr) [See also Figure 8 ($m_{1}=3M_{\odot}$) and Figure 9 ($m_{1}=20M_{\odot}$) in Appendix A, which demonstrates S/Ns as a function of observation time $T_{\rm obs}$ for different $m_{1}$ and detectors.]. If the primary object is an astrophysical black hole, its mass should be greater than $3M_{\odot}$, placing it in the region to the right of the black vertical dotted line in Figure 2 and 3.
@@ II.5 Uncertainty of Parameters
In GW data analysis, we usually utilize Fisher information matrix to estimation the error or uncetrainty of parameters (2020MNRAS.496..182L; 2023MNRAS.522.2951Z; 2025arXiv251213151Z). For two parameters $p_{\alpha}$, $p_{\beta}$ of GW system, the Fisher information matrix (FIM) is defined as follows:
@@
where $|\tilde{h}(f)|\propto\mathcal{M}^{5/6}$ is the waveform of its GW during its inspiral stage in frequency domain (1996PhRvD..53.2878F; 2018MNRAS.476.2220L). Thus it is evident that —∂~h∂M—=56—~h—M. Substitute it into Equation (10), finally we obtain
@@ III Eccentric Orbit
2025arXiv251213151Z; 2025ApJ...995...27X demonstrate that binary systems residing in globular clusters or nuclear clusters can retain nonzero orbital eccentricity.
If the orbit of a binary system is not circular but eccentric, the frequencies of the GWs it radiates will no longer be simply twice the orbital frequency $f_{\rm orb}$. Instead, there will exist many higher-order harmonic frequencies at integer multiples of the orbital frequency $nf_{\rm orb}$, n=1,2,3,n=1,2,3,...(1963PhRv..131..435P; maggiore2008gravitational). Such a GW signal is easier to analyze in the frequency domain.

### v2

@@ Abstract
The recent discovery of M31 UCXB-1, the first extragalactic ultracompact X-ray binary (UCXB) with an orbital period of $T_{\rm orb}\sim 465$ s, presents a unique laboratory for studying close binary evolution and an unprecedented target for continuous gravitational wave (GW) searches. Its identification as a strong candidate black hole-white dwarf (BH-WD) system, combined with its exceptionally short period and high X-ray luminosity, suggests it may be one of the most vital low-frequency GW sources in M31. In this paper, we investigate the detectability of its GW signal for future space-borne detectors in multiband GW detection. We find that while its signal-to-noise ratio (S/N) for low-frequency detectors remains marginal for high-confidence detection, middle-frequency detectors such as DECIGO and BBO are far more promising, potentially achieving S/N $\varrho>8$ within reasonable observational duration. With a primary mass of only $m_{1}>5.4M_{\odot}$ (or $6.6M_{\odot}$), the network of all low and middle frequency detector (or BBO alone) is capable of detecting GW from this system with a $\varrho>8$, during 10-year observation. Furthermore, orbital eccentricity can enhance the GW strain at higher harmonics, further improving its detectability, especially for middle-frequency detectors. This study establishes M31 UCXB-1 as a key prototype of short-period UCXBs, cementing its role as a cornerstone for multiband, multi-messenger astrophysics and a vital bridge between X-ray astronomy and the future GW era.
@@ I Introduction
Ultracompact X-ray binaries (UCXBs), with orbital periods below $\sim 80$ minutes, represent the end state of low-mass X-ray binary evolution involving a compact accretor [neutron star (NS) or black hole (BH)] and a hydrogen-deficient donor (Wang et al., 2021; Tauris and van den Heuvel, 2023). They are powerful laboratories for studying extreme accretion physics, tidal interactions, and angular momentum loss mechanisms (e.g., Chen and Podsiadlowski, 2016; Chen et al., 2021). Furthermore, their tight orbits make them prominent sources of persistent, low-frequency gravitational waves (GWs) driven by radiation reaction, placing them squarely within the sensitivity bands of proposed space-borne interferometers (Nelemans, 2009; Chen et al., 2020; Chen, 2020; Qin et al., 2023; Suvorov, 2021; He et al., 2023; Xuan et al., 2025). including Laser Interferometer Space Antenna (LISA, Amaro-Seoane et al. 2017), Tianqin (Luo et al., 2016), and Taiji (Huang et al., 2017).
@@
Recently, a remarkable system has emerged from deep X-ray monitoring of the Andromeda galaxy (M31). Zhang et al. (2024) reported the discovery of a periodic X-ray source (RX J0042.3+4115/3XMM J004222.9+411535) in the M31 bulge with a period of $T_{\rm orb}\approx 465$ s and a peak luminosity of $L_{X}\sim 2.9\times 10^{38}$ erg s-1. Follow-up spectral analysis by Ma et al. (2026) strongly argues that this source, designated M31 UCXB-1, is the first known extragalactic UCXB and a compelling BH-white dwarf (WD) candidate. It possesses the shortest orbital period among known UCXBs. The unique parameters of M31 UCXB-1—a likely BH-WD binary with a $\sim 10^{2}$-s period at a distance of 785 kpc—make its GW emission exceptionally interesting. The expected GW frequency ($f_{\rm gw}\approx 4.3$ mHz) falls within the sensitivity range of low-frequency GW observatories. Previous estimates suggest it may be detectable by LISA/Taiji if the BH mass is sufficiently high (Yang et al., 2025; Ma et al., 2026). However, a detailed and systematic assessment of its detectability across the full landscape of future GW missions, considering middle-frequency detectors, mission durations, parameter estimation error and the potential impact of orbital eccentricity, is still lacking.
@@
In this paper, we perform a comprehensive investigation into the GW detectability of M31 UCXB-1 in low and middle frequency band. We go beyond simple characteristic strain comparisons and compute expected signal-to-noise ratios (S/N) for a suite of planned low and middle frequency detectors, including LISA, Taiji, TianQin, DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) (Kawamura et al., 2006), and Big Bang Observer (BBO) (Harry et al., 2006). We explore a realistic range of system parameters (BH mass, orbital eccentricity) informed by the latest observational and theoretical constraints (Zhang et al., 2024; Ma et al., 2026; Yang et al., 2025). Our analysis reveals that middle-frequency detectors offer a superior pathway for a high-confidence detection of this source. Furthermore, we quantify that if the system exhibits non-zero orbital eccentricity, it can enhance the detection S/N for both low- and middle-frequency GW detectors, with a particularly pronounced improvement for middle-frequency detectors. This work solidifies M31 UCXB-1’s role as a prime multimessenger and multiband target and provides critical predictions for the observing strategies of forthcoming GW observatories in low and middle frequency band.
@@
This paper is structured as follows: In Section II, we assess the GW detectability of M31 UCXB-1 and the associated parameter estimation uncertainties for low- and middle-frequency GW detectors, under the assumption of a circular orbit. Subsequently, in Section III, we extend the analysis to include eccentric orbital configurations. Finally, we summarize our findings and present conclusions in Section IV.
@@ II.1 Parameter Setting
The binary system is located at a distance of approximately 785 kpc from Earth. Its orbital period is about 465 s, corresponding to an orbital frequency of $2.15\times 10^{-3}$ Hz. The WD in the system has a mass of $0.08761\,M_{\odot}$, while the mass of the black hole remains uncertain due to its degeneracy with the orbital inclination in X-ray observations (Ma et al., 2026). A BH mass of $m_{1}=3M_{\odot}$ corresponds to an inclination $\iota\geq 73.5^{\circ}$, while $m_{1}=20M_{\odot}$ corresponds to $\iota\geq 86.7^{\circ}$ (Ma et al., 2026). Since the amplitude of GW is not sensitive to the inclination $\iota$ (see Table 2), and eclipsing phenomena reflects this system is basically edge on. Unless otherwise specified, throughout this work we adopt an edge-on orbital inclination of $\iota=90^{\circ}$. Some fixed parameters about the system are summarized in Table 1. Since Ma et al. (2026) assume that its orbit is circular, we just consider circular orbit in this Section, and we will consider its eccentric orbit in next Section (Sec. III).
@@ II.2 GW Signals
The GW radiated from a binary system can be expressed as (Maggiore, 2008; Guo et al., 2025)
@@
where $h_{0}=\frac{4{G}^{5/3}\mathcal{M}^{5/3}(\pi f)^{2/3}}{{c}^{4}d}$, $f=2f_{\rm orb}$ is the frequency of GW, $A_{+}=\frac{1+\cos^{2}\iota}{2}$, $A_{\times}=\cos\iota$, $\times$ represent different polarizations, $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass of this system and $\Phi(t)$ is the phase of GW. The root-mean-square (RMS) average of GW strain on all polarization is h_rms=⟨h_+^2+h_×^2⟩=h_0A+2+A×22. Adopting matched filtering method, the characteristic strain for GW is (see Moore et al., 2015a; Sesana et al., 2005)
@@
where $N_{\rm eff}=\min\left\{\frac{f^{2}}{\dot{f}},T_{\rm obs}f\right\}$(Guo et al., 2022, 2025), the variation rate of frequency is $\dot{f}=\frac{96}{5}\pi^{8/3}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{5/3}f^{11/3}$. For this binary system, the relation $\frac{f^{2}}{\dot{f}}\gg T_{\mathrm{obs}}f$ holds, implying that the evolution timescale of the GW frequency is much longer than the coherent observing duration. Consequently, the characteristic strain $h_{\mathrm{c}}$ depends primarily on the observation duration $T_{\mathrm{obs}}$.
@@
And its merger time scale due to GW radiation can be estimated by (Maggiore, 2008)
@@
Therefore, we can summarize several typical characteristic strains and other dynamical parameters for different cases in Table 2. Although the accretion process of a black hole can alter the masses $m_{1}$ and $m_{2}$ of the binary system, the resulting mass change is negligible over the observational timescale (<10<10 yr) due to low accretion rate $\dot{M}\sim 10^{-8}M_{\odot}$/yr (Yang et al., 2025)]. Therefore, in our study on GWs, we neglect the mass variation due to accretion as well as the potential changes in orbital dynamics.
@@ II.3 GW Detectors
Tianqin $10^{-4}$ 0.5 Liu et al. (2020b)
LISA $10^{-4}$ 0.5 Robson et al. (2019)
Taiji $10^{-4}$ 0.5 Liu et al. (2023)
DECIGO $10^{-3}$ $10^{2}$ Yagi and Seto (2011, 2017)
BBO $10^{-3}$ $10^{2}$ Yagi and Seto (2011, 2017)
It is likely that this system will be detected by future low-frequency and middle-frequency GW detectors. The low-frequency GW detectors encompass LISA, Taiji, and Tianqin, whereas the middle-frequency GW detectors include DECIGO and BBO in this paper. For all GW detectors used in this work, their noise power spectral density $S_{\rm n}(f)$ are summarized in the references provided in Table 3 and their detectable frequency ranges are also summarized in Table 3. We show some typical characteristic strain values of this GW source (red and green points/diamonds, Eq. 4) and noise curve of all kinds of GW detectors in low and middle frequency band in the Figure 1. The characteristic strain values computed by us are broadly consistent with those reported by Yang et al. (2025); Ma et al. (2026). The slightly lower amplitude in our results is due to our choice of orbital inclination $\iota=90^{\circ}$.
@@
Figure 1: Possible characteristic strain $h_{\rm c}(f)$ of GW emitted from this system [red ($m_{1}=3M_{\odot}$) and green ($m_{1}=20M_{\odot}$) points($T_{\rm obs}=4$ yr)/diamond($T_{\rm obs}=10$ yr)] and noise curve $h_{\rm n}(f)$ for different GW detectors in low and middle frequency band with different colors represent different GW detectors as shown in the legend. These dotted lines for Taiji and LISA represent the noise curves of GW detectors without Galactic binaries foreground confusion (See Liu et al., 2023). And this black line with triangle (red line with $\times$) symbols represents characteristic strain $h_{{\rm c},i}$ for eccentric orbit with e=0.2e=0.2(0.4), $m_{1}=3M_{\odot}$, and $T_{\rm obs}=4$ yr.
@@ II.4 S/N
The S/N of GW from this system detected by a GW detector is (Moore et al., 2015a, b)
@@
For a network of GW detectors, its effective S/N is calculated as follows(Guo et al., 2022, 2025)
@@
where $\varrho_{i}$ is the S/N for each detector $i$, the sum is taken over the $N_{\rm d}$ detectors. In this paper, we mainly consider two GW detector networks: Tianqin-LISA-Taiji low frequency GW detector network and Tianqin-LISA-Taiji-DECIGO-BBO (TLTDB) network including all low and middle frequency GW detectors in this paper. Through calculation, we obtain its S/Ns for different detectors as a function of $m_{1}$, as shown in Figure 2 ($T_{\rm obs}=4$ yr) and Figure 3 ($T_{\rm obs}=10$ yr) [See also Figure 8 ($m_{1}=3M_{\odot}$) and Figure 9 ($m_{1}=20M_{\odot}$) in Appendix A, which demonstrates S/Ns as a function of observation time $T_{\rm obs}$ for different $m_{1}$ and detectors.]. If the primary object is an astrophysical black hole, its mass should be greater than $3M_{\odot}$, placing it in the region to the right of the black vertical dotted line in Figure 2 and 3.
@@ II.5 Uncertainty of Parameters
In GW data analysis, we usually utilize Fisher information matrix to estimation the error or uncetrainty of parameters (Liu et al., 2020a; Zhao et al., 2023, 2025). For two parameters $p_{\alpha}$, $p_{\beta}$ of GW system, the Fisher information matrix (FIM) is defined as follows:
@@
where $|\tilde{h}(f)|\propto\mathcal{M}^{5/6}$ is the waveform of its GW during its inspiral stage in frequency domain (Finn, 1996; Li et al., 2018). Thus it is evident that —∂~h∂M—=56—~h—M. Substitute it into Equation (10), finally we obtain
@@ III Eccentric Orbit
Zhao et al. (2025); Xuan et al. (2025) demonstrate that binary systems residing in globular clusters or nuclear clusters can retain nonzero orbital eccentricity.
If the orbit of a binary system is not circular but eccentric, the frequencies of the GWs it radiates will no longer be simply twice the orbital frequency $f_{\rm orb}$. Instead, there will exist many higher-order harmonic frequencies at integer multiples of the orbital frequency $nf_{\rm orb}$, n=1,2,3,n=1,2,3,...(Peters and Mathews, 1963; Maggiore, 2008). Such a GW signal is easier to analyze in the frequency domain.